### v2

# Entropy-Memorization Law: Evaluating Memorization Difficulty of Data in LLMs

Abstract
Large Language Models (LLMs) are known to memorize portions of their training data, sometimes reproducing content verbatim when prompted appropriately. In this work, we investigate a fundamental yet under-explored question in the domain of memorization: How to characterize memorization difficulty of training data in LLMs? Through empirical experiments on OLMo, a family of open models, we present the Entropy–Memorization Law. It suggests that data entropy is linearly correlated with memorization score. Adopting the same strategy to discover Entropy-Memorization Law, we derive a simple yet effective approach to distinguish training and testing data, enabling Dataset Inference (DI).
1 Introduction
Language modeling has become a central task in modern natural language processing (NLP). Large Language Models (LLMs) demonstrate remarkable performance in capturing linguistic patterns and generating coherent text [1, 2, 3]. Alongside dominating performance through various benchmarks, a critical phenomenon has emerged: LLMs are shown to memorize and reproduce verbatim sequences from their training corpora [4, 5]. This memorization behavior has raised growing concerns, particularly in terms of privacy leakage and intellectual property protection. For example, studies have shown that LLMs can inadvertently generate personally identifiable information (PII) [5], or proprietary data from books [6, 7] and news articles [8].
As scaling laws [9] drive LLM developers to expand model capacity and training data for performance improvements, research [10, 11] has demonstrated that memorization scales with model size. Broader data exposure in LLM training elevates the risk of leakage for all internet-sourced content. Furthermore, memorization is shown to be necessary for generalization [12].
Therefore, advancing the theoretical understanding of the factors that shape memorization has become a crucial and urgent issue in LLM development. Previous studies explore several factors, including prompting strategy [5, 13], and training paradigm [14]. This study aims to address an open and under-explored question: How to characterize memorization difficulty of training data in LLMs? So far, the few preliminary explorations through this question focus on deduplication – duplicated data significantly increases memorization [15, 16, 10]. Exploring this question facilitates the assessment of privacy risks for LLM providers. The manifestation of memorization is typically measured by a memorization score. To characterize “memorization difficulty”, this work fixes other factors in memorization and looks for an approximator function of memorization score. This function takes input from training data, and aims to approximate the memorization score of the data.
Using heuristics from information theory to build the approximator function, we present the core finding of this study: Entropy–Memorization Law – entropy estimator approximates memorization score well. It suggests that higher entropy correlates with higher memorization scores in LLMs (with $r>0.90$). This law is empirically verified on the OLMo family [17] and its accompanying open datasets. Moreover, we find that in EM Law, (i) lower-entropy data comprises fewer tokens, and achieves near-maximal entropy values, while (ii) higher-entropy data contains exponentially more tokens but exhibits highly non-uniform distributions compared to lower-entropy data. Furthermore, by slightly modifying the strategy to discover the Entropy-Memorization Law, we derive a simple yet effective approach to distinguish training and testing data, enabling Dataset Inference (DI). In practice, DI could be used to detect test data contamination and unauthorized use of proprietary data.
2 Experimental Setup
Threat Model
This paper assumes a hypothetical engineer who studies the characterization of training data on an LLM. Therefore, it is necessary for the engineer to have full access to the LLM and its training data.
Beyond the training data, other co-founders can shape the memorization score. Co-founders include prompt strategy, training paradigm (we will discuss in detail in Sec. 7). This paper then fixes all the possible co-founders, which will be explained in the following.
LLM and Its Training Corpus
This paper employs pre-trained-only LLM, since post-training may involve different model training paradigms, bringing uncontrollable noises to our experiments. We select the OLMo family, a fully open LLM family. It achieves state-of-the-art performance compared with other fully-open LLMs like Pythia [16] and DCLM [18]. OLMo-series comes with WIMBD [19], an open-source tool to perform efficient data count and search on their training set.
We selected two pre-trained only LLMs: OLMo-1B [17] pre-trained on Dolmo [20] dataset, and OLMo-2-1124-7B [21] pre-trained on OLMo-2-1124-Mix [21] dataset. The datasets were deduplicated to address the well-documented phenomenon where data deduplication significantly increases memorization. All datasets are licensed with Open Data Commons Attribution License (ODC-By).
Prompting Strategies
In this study, we employ Discoverable Memorization (DM) [22, 23, 15, 11]. Formally, DM denotes the following strategy: we sample $N$ token sequences from the training dataset. Each sequence is partitioned into $(p,s)$, where $p$ (first $|p|$ tokens) serves as the prompt and $s$ (remaining tokens) serves as the answer for model $\theta$. Afterwards, LLM $\theta$ generates response $r=\theta(p)$. The memorization score measures the difference between two sequences $r$ and $s$.
The token sequence sampling strategy is as follows: we repeatedly randomly sample a sequence (length > $|p+s|$) from the dataset until the number reaches the required number.
In this experiment, we set $|p|=100,|s|=|r|=50$ unless otherwise specified, following the popular setup in the community [24]. We sampled N=240, [^0] N=240{,}000 times on Dolma, and N=300, [^0] N=300{,}000 times on OLMo-2-1124-Mix. In the supplementary materials, we include the detailed statistics of the sampled dataset, and experimental results with LLM generating in varying lengths.
Filtering Trivial Memorization
Empirical observations show that LLM may trivially memorize – the LLM response $r$ may exhibit a high degree of lexical overlap with the prompt $p$. For instance, LLM may copy a long URL from the prompt as the response. Although such a response may match the answer, however, such a match should be attributed to the prompt, instead of the memorization capacity of LLM. Therefore, filtering out cases of trivial memorization helps to calibrate our analysis on memorization. To conduct filtering, we design a Longest Common Subsequence (LCS)-based filtering method. We empirically establish a thresholding strategy: samples for which $LCS(p,s)\geq\frac{|s|}{2}$ are excluded from further memorization analysis, while samples below the threshold are retained.
Memorization Score
In our context, a memorization score $d(r,s)$ measures the differences between response $r$ and answer $s$ at the token level. Following previous memorization work [25], we use Levenshtein Distance, also known as edit distance [26]. It is defined by the minimal number of single-token edit operations – insertions, deletions, and substitutions – required to transform one sequence into another. A higher memorization score indicates lower similarity between two sequences.
Depending on the notions of memorization, other memorization scores are adopted in the community. The traditional one is exact match [23, 27, 4, 5], a binary decision on whether two sequences match exactly or not. Such a score ignores the partial memorization, which can be crucial for assessing privacy risks. Hence, it is undesirable in our work. Another line of research uses semantic similarity, measuring the similarities based on sentence embeddings generated by LMs [28, 29]. However, memorization at the semantic level could be subject to human decisions. Therefore, Levenshtein distance is a good fit as a memorization score in our study.
To summarize, the research question in this paper is formulated as follows:
Assumption. A fixed pre-trained LLM $\theta$, a fixed prompting strategy $DM$ to generate $p$, and a memorization score $d(r,s)=d(\theta(p),s)$. Goal of the study. Find an approximator function $M(s)$ of memorization score $d(r,s)$.
3 The First Attempt: An Instance-Wise Entropy Estimator
How to characterize training data $s$? A natural idea is to use entropy  [^1] We assume a base-2 logarithm for all entropy calculations throughout the work. – we suspect that higher-entropy answers exhibit greater randomness, and may be harder to be memorized. Therefore, in our first attempt, we characterize $M(s)$ using a histogram-based entropy estimator [30]. The entropy is estimated per instance. To clear any ambiguities, we use the formal notations to describe the calculation:
In our context, an instance $s_{i}=(s_{i}^{1},s_{i}^{2},...,s_{i}^{|s_{i}|})$ is a sequence, where $s_{i}^{j}$ is a token. All tokens within $s_{i}$ form the sample space $\mathcal{T}_{i}$. Then for each token $x\in\mathcal{T}_{i}$, the empirical point probabilities $\hat{p}_{i}(x)$ are calculated as:
$\hat{p}_{i}(x)=\frac{1}{|s|}\left|\{j\mid s_{i}^{j}=x\}\right|.$ (1)
$\hat{p}_{i}(x)$ is the relative frequency of $x$ in the observed sequence. In this attempt, we use entropy estimated by the empirical point probabilities as our approximator $M(s_{i})$:
$M(s_{i})\triangleq-\sum_{x\in\mathcal{T}_{i}}\hat{p}_{i}(x)\log\hat{p}_{i}(x)$ (2)
With the established $M(s_{i})$, we are interested in whether $M(s_{i})$ is a good approximator of the memorization score $d(r_{i},s_{i})$. To achieve this, we gather all $(M(s_{i}),d(r_{i},s_{i}))$ pairs obtained by empirical observations in a scatter plot, and further study their correlation. The detailed algorithm is as follows:
Input: LLM $\theta$, and its training corpus $D$
Output: Plot of $(d(r_{i},s_{i}),M(s_{i}))$
[…]
2for $i\leftarrow 0$ to N1N-1 do
3 $r_{i}\leftarrow\theta(p_{i})$;
 $\hat{p}_{i}\leftarrow$ EmpProb(r,s)(r,s)// Eq. 1
   $M(s_{i})\leftarrow-\sum_{x\in\mathcal{T}_{i}}\hat{p}_{i}(x)\log\hat{p}_{i}(x)$// Eq. 2
4   $d(r_{i},s_{i})\leftarrow d_{\text{lev}}(r_{i},s_{i})$
5  Plot $(d(r_{i},s_{i}),M(s_{i}))$;
6
7 end for
Algorithm 1 Instance-wise entropy estimator.
Figure 1: Memorization score v.s. entropy estimator observed on OLMo-1B.
We run algorithm 1 on the OLMo-1B model, and obtain the scatter plot as illustrated in Fig. 1. A linear regression based on ordinary least squares [31] is conducted on the plot.
The regression analysis suggests a positive linear relationship between Levenshtein Distance-based memorization score and our approximator. However, the observed pattern is very noisy, as demonstrated by the weak Pearson correlation $r=0.581$.
We regard the failure of our first attempt as due to the limitation of sample space. In our experimental setup, an LLM generates up to $|s_{i}|=50$ tokens. Hence, the sample space size $|\mathcal{T}_{i}|$ is upper bounded by 50 in ideal cases. This space is orders of magnitude smaller than the full token space. In reality, the full token space is defined by the tokenizer vocabulary, and the vocabulary size of the OLMo-1B tokenizer is about 50k. Therefore, our entropy estimation is too noisy to reflect the real-world scenarios. Our second attempt addresses the limitation and will be discussed in the next section.
4 The Second Attempt: A Level-Set-Based Entropy Estimator
Our first attempt’s failure is attributed to insufficient sample space for entropy estimation. The second attempt addresses this by substantially expanding the sample space. We observe that in the first attempt (Fig. 1), for a fixed memorization score, we constructed more than one (even thousands of) estimates. Then an intuitive idea is to aggregate these estimates to a new robust estimate. Specifically, we expand the sample space from tokens with one instance, to all the instances with the same memorization score. The idea is the same as building “level-sets” of memorization score. Mathematically, for a fixed memorization score $e$, the new sample space
$\mathcal{T}_{e}=\bigcup\left\{s_{i}^{j}\mid d(r_{i},s_{i})=e,i\in\{0,\cdots,N-1\},j\in\{0,\cdots,|s|-1\}\right\}.$ (3)
Then the empirical probabilities $\hat{p}_{e}(x)$ are calculated within new sample space $\mathcal{T}_{e}$:
$\hat{p}_{e}(x)=\frac{1}{N|s|}\left|\{(i,j)\mid s_{i}^{j}=x,d(r_{i},s_{i})=e\}\right|.$ (4)
Input: LLM $\theta$, and its training corpus $D$
Output: Plot of $\left(e,M(s_{e})\right)$
[…]
2for $i\leftarrow 0$ to N1N-1 do
3  $r_{i}\leftarrow\theta(p_{i})$;
4  $d(r_{i},s_{i})\leftarrow d_{\text{lev}}(r_{i},s_{i})$
5 end for
6for $e\leftarrow 0$ to |s|1|s|-1 do
  $\hat{p}_{e}\leftarrow$ NewEmpProb($r,s$) // Eq. 4
7  $M(s_{e})\leftarrow-\sum_{x\in\mathcal{T}_{e}}\hat{p}_{e}(x)\log\hat{p}_{e}(x)$;
8  Plot $\left(e,M(s_{e})\right)$.
9 end for
10
Algorithm 2 Level-set-wise entropy estimator.
The new empirical probabilities benefit from a larger sample space. Similar to Eq. 2, we then use new empirical probabilities to derive a new level-set-based entropy estimate to approximate the memorization score $e$.
$M(s_{e})\triangleq-\sum_{x\in\mathcal{T}_{e}}\hat{p}_{e}(x)\log\hat{p}_{e}(x)$ (5)
The modified algorithm is as shown in Alg. 2. It turns out that the level-set-based entropy estimator is shown to be a good approximator of memorization score, as demonstrated below.
5 Entropy-Memorization Law
We run algorithm 2 on OLMo-1B and OLMo-2-1124-7B, and the empirical results are shown in Fig. 2(a) and  2(b). We observe very strong linear empirical results ($r=0.972$ and $0.945$ respectively) on both plots. It indicates that the level-set-based entropy estimator is an effective linear approximation of memorization score. We name this discovery as Entropy-Memorization Law, or EM Law. Additionally, we made the following observations:
  • Larger models show steeper slope, and lower intercept in the regression.
  • The zero-distance point, i.e., $(0,M(s_{0}))$ exhibits a significant deviation from the regression line in both plots.
  • Although two regression lines have different slopes and intercepts, the $y$-value reaches about 11 when the memorization score $e=50$. Besides, $M(s_{50})\approx 11$.
(a) OLMo-1B
(b) OLMo-2-1124-7B
Figure 2: Memorization score v.s. Level-set-wise Entropy Estimator
5.1 Inside Entropy–Memory Law
Instead of finding an even better approximator, this subsection studies the properties of the approximator that we obtained in our second attempt to facilitate a deeper understanding.
In our framework, entropy is determined by two key factors: (1) the cardinality of the possible outcomes (sample space size), and (2) the distribution of empirical probabilities across these outcomes. Mathematically, given $n$ possible outcomes, the entropy of a discrete random variable (r.v.) is upper bounded by $\sum_{i=1}^{n}\frac{1}{n}\log\frac{1}{n}=\log n$. The upper bound is achieved when r.v. is uniformly distributed. Hence, a two-outcome r.v. has maximum entropy $\log 2=1$ bit, while a three-outcome r.v. yields maximum $\log 3$ bit entropy. So far, we have developed an entropy estimator. Then, we are interested in the roles of these two factors in entropy calculation.
For each memorization score, we report the sample space size formed in Fig. 3. The statistics are calculated following the same experimental setup on OLMo-2-1124-7B. Specifically, we depict $(e,\mathcal{T}_{e})$ pairs using green cross sign $\times$. The experimental result indicates that the sample space, or unique token count, grows exponentially as memorization score increases. Remarkably, while the full dataset contains tens of millions of tokens, perfect memorization (Levenshtein distance=0) occurs within merely $2^{6}$ unique tokens.
Figure 3: Memorization score v.s. Normalized entropy on OLMo-2-1124-7B. Note: “unique token count” scale is exponential.
After obtaining the sample space size, we examine how empirical probabilities are distributed in the space. It is characterized by normalized entropy:
$\overline{M}(s_{e})\triangleq\frac{M(s_{e})}{H_{\max,e}}=\frac{M(s_{e})}{\log|\mathcal{T}_{e}|},$ (6)
where $\overline{M}(s_{e})$ normalizes the entropy estimate $M(s_{e})$ by its theoretical maximum $H_{\max,e}$. Values approaching 1 indicate a near-uniform distribution over the token set $\mathcal{T}_{e}$, while lower values suggest greater non-uniformity.
Following the same setup on OLMo-2-1124-7B, we plot $(e,\overline{M}(s_{e}))$ using blue dots on Figure 3 to observe how the normalized entropy estimator changes with the memorization score. Through normalized entropy, we observe that the empirical distributions maintain above 0.85 for memorization scores ranging from 0 to 40. We observe a trend that normalized entropy decreases as the memorization score increases.
To summarize, our findings indicate that (i) lower-entropy data, comprises fewer tokens, and achieves near-maximal entropy values, while (ii) higher-entropy data contains exponentially more tokens but exhibits highly non-uniform distributions compared to low-entropy cases.
6 Entropy-Memorization Law on Test Data
Running algorithm 2 on training dataset of LLMs, we discovered Entropy-Memorization Law. This section then explores another question that naturally arises – what happens if we run the same algorithm on test dataset? It turns out that the plot behaves very differently from training datasets. Detailed experimental results will be presented later in this section. We thus derive a simple strategy to tell the difference between train data and test data for LLM. In the AI community, this task is studied under the name Dataset Inference. The following content describes the background of DI, and presents experimental results on DI tasks.
Dataset Inferences
Dataset Inference (DI) [37, 38] builds on the idea of membership inference attacks (MIA). While both MIA and DI aim to identify whether some suspect data was part of the training data, they differ on the amount data required. MIA operates at instance (sentence) level, however, DI operates on a collection of instances – in reality, the suspect data used for DI could be a book.
In practice, DI can identify potential test set contamination, to provide a calibrated performance evaluation of LLMs. DI may also detect unauthorized usage of copyrighted training data, thus promoting the protection of intellectual properties. Moreover, by using more data to determine membership, DI is deemed to be more realistic than MIAs. As Maini et al. [38] presented that as the size of the training set increases, the success of membership inference degrades to random chance.
Experimental Results
We run algorithm 2 slightly abusing the notation on multiple LLMs and datasets. A linear regression analysis is further conducted, and the corresponding intercept and slope are reported.
Selected LLMs. We select OLMo-2-1124-7B (“OLMo-2”), Pythia-6.9b-deduped [16] (“Pythia”).
Selected Datasets. We select LiveBench [39] and MIMIR [40]. For LiveBench, we use data from 2024-06-25 to 2024-11-25. The data is after the release of GPT-2-XL and OLMo-2-1124-7B, hence it is guaranteed to be non-member data (label 0). MIMIR is a dataset originally for evaluating MIAs on the Pythia suite by re-compiling the Pile [41] train/test splits. We use “Pile CC” subset for evaluation.
Table 1: Dataset Inference Results. “Pred.” denotes the prediction of our method EMBEDI.
Model Dataset Intercept Slope Pred. Label
OLMo-2 LiveBench 2.202 0.155 0 0
GPT-2-XL LiveBench 1.148 0.159 0 0
Pythia MIMIR_cc -2.048 0.251 0 0
Pythia MIMIR_cc 3.992 0.091 1 1
Figure 4: Visualization of intercepts and slopes in DI.
The experimental results are summarized in Table 6 and Figure 4. We empirically observe that the plots of training data and testing data exhibits a disparate regression line. Dataset membership could be determined by simply thresholding intercepts or slopes resulted from algorithm 2. Member-data can be characterized by slope $m>\tau_{m}$, or intercept $k>\tau_{k}$, while non-member data is characterized by slope $m\leq\tau_{m}$, or intercept $k\leq\tau_{k}$. We dub this threshold-based strategy as EMBEDI (Entropy-Memorization law-BasEd Dataset Inference). It is a reference-free strategy, without any training required. For more discussions on the experimental results, we refer the readers to supplementary materials.
7 Related Work
7.1 Memorization
Since the discovery of the memorization phenomenon in the late 2010s [42, 4, 5, 43], the AI Security and Privacy research community has maintained a strong interest in the phenomenon and its implications. The following paragraphs examine how memorization in language models is influenced by key factors, including training data, model paradigm, and prompting strategy.
Data shapes memorization. Several literature suggests that [15, 16] duplicated data significantly increases memorization. Larger models trained on bigger datasets show increased memorization [44, 16]. Other studies [27, 10] investigate how memorization manifests across data with varying semantics and sources.
Model Paradigm shapes memorization. Beyond pre-trained language models, recent work has explored memorization in post-training stages. Chu et al. [14] demonstrate that supervised fine-tuned (SFT) LLMs exhibit stronger memorization tendencies than those trained with reinforcement learning (RL). Additionally, Nasr et al. [22] reveal that safety-aligned models still retain memorized data.
Prompting Strategy shapes memorization. Researchers employ three main types of prompting strategies for language models, categorized by threat models. A significant body of work relies on manual efforts or template-based approaches to generate prompts at scale, as seen in [4, 5, 45]. Studies such as [5, 15, 46] demonstrate that longer prompts substantially increase the likelihood of reproducing memorized training data sequences. Another line of research constructs prompts directly from existing data sources, such as training corpora or web data [22, 23, 15, 11, 47]. Recent advances involve more sophisticated strategies that leverage synergies between LLMs and training data. For instance, Zhang et al. [48] quantify how a model’s performance on an example $x$ depends on whether $x$ was included in the training data. Additionally, [13] adapts GCG [49]—a prompt optimization tool originally designed for adversarial attacks—to generate effective extraction prompts.
7.2 Dataset Inference
Aiming to determine whether an instance is included in the LLM training corpus or not, Membership Inference Attacks (MIA) [50, 51, 52, 53, 54, 55, 56, 57, 58] are very popular in the AI Privacy community, due to their promising applications to detect uncopyrighted data use in LLM training. Therefore, MIA was also studied under other names like “detect test set contamination”, or ”detect pre-training data” [59, 60, 61]. There are hot debates within the community that MIA learns shortcuts hence exhibits limited robustness in realistic settings [40, 62]. In contrast, Dataset inference [38, 37, 63] is less popular, but may have the potential to exhibit higher robustness since it requires a large amount of data.
8 Limitations and Broader Impacts
8.1 Limitations
Empirical experiments. Our work adopts a single set of prompting strategy (DM) and memorization score (edit distance) in our memorization experiments. Although the setup is commonly used by other studies, other combinations exist. We would like to explore adversarial compression [13], and non-adversarial reproduction [47] in our future work.
Under-explored theoretical explanations of EM Law. Using heuristics in information theory, this work finds a linear approximator of the memorization score. However, why the proposed level-set-based entropy estimator works remains underexplored. Potential theoretical tools include the long-tail theory by Feldman and other researchers [43, 12], and multi-calibration in LLMs [64].
Memorization score prediction. Due to limitation of sample space, our strategy does not enable memorization score prediction, nor “difficulty prediction” at the instance level.
8.2 Implications and Societal Impact
This work facilitates the theoretical understanding of factors in LLM memorization. Beyond that, we made a few observations that might be impactful to society:
  • For trainers of LLMs, our work contributes a guideline in model audition: lower entropy data is at higher risk of leakage.
  • For online service providers, our study finds a potential defect in credential design. Specifically, maximizing char-level entropy does not guarantee minimum chances of credential leakage from LLMs. Instead, some credentials are easiest to memorize.
  • We develop a simple yet effective strategy to detect the use of proprietary data in LLMs. It facilitates the protection of intellectual property. Such a strategy could be adopted to detect test set contamination, enabling a calibrated performance evaluation of LLMs.
To the best of the authors’ knowledge, this research does not introduce any negative societal impacts.
9 Takeaways
Here are the summarized results of this paper:
  1. 1.
    We present the Entropy-Memorization Law: high entropy training data linearly correlates with a high memorization score. The law is discovered when we approximate the memorization score using a level-set-based entropy estimator.
  2. 2.
    A further investigation indicates that in EM Law, (i) lower-entropy data comprises fewer tokens, and achieves near-maximal entropy values, while (ii) higher-entropy data contains exponentially more tokens but exhibits highly non-uniform distributions compared to lower-entropy data.
  3. 3.
    By slightly modifying the strategy to discover the Entropy-Memorization Law, we derive a simple yet effective approach to distinguish training and testing data. It can serve as a baseline method for Dataset Inference (DI), a variant of Membership Inference Attacks.
Appendix A Compute
All experiments were conducted on a GPU cluster equipped with 4 NVIDIA RTX 3090 GPUs (24GB CUDA memory per card), running Ubuntu 22.04. We use VLLM [65] to speed up inference.
Appendix B Details on Experimental Setup
B.1 Composition of Sampled Dataset
We constructed datasets with sizes 240,000 and 300,000, respectively, for Dolma and OLMo-2-1124-Mix. Table 3 and 3 presents the composition of sampled datasets after LCS filtering.
Table 2: Source Counts of Dolma
Source Dataset Count
pes2o 39,777
$c$ 39,743
books 39,598
reddit 39,373
stack 38,865
wiki 23,115
Total 219,186
Table 3: Source Counts of OLMo-2-1124-Mix
Dataset Count
algebraic-stack 797
arxiv 1,464
dclm1 28,130
dclm2 27,984
dclm3 28,014
dclm4 28,163
dclm5 28,049
dclm6 28,065
dclm7 28,060
dclm8 28,091
dclm9 28,040
dclm10 28,032
open-web-math 499
pes2o 4,364
starcoder 5,280
wiki 276
Total 293,308
Appendix C Extended Results on Entropy-Memorization Law
C.1 Sampling Strategy of LLMs
In the main body of the paper, we assume a fixed temperature of 0.8. In this subsection, we adopt different sampling strategies of LLMs and discuss how these strategies might shape EM Law. Due to computation constraints, we conduct our experiments on a subset “DCLM1” with OLMo-2-1124-7B. The size of the subset is around 28,000.
We consider combinations of temperature, Top-k sampling, and nucleus sampling (top-p). The experimental results are summarized in Tab. 4, and details are shown in Fig. 5 -  10. Under all sampling strategies we have explored, we empirically observe that EM Law holds with $r>0.92$. Beyond that, we made a few observations here:
  • The zero-distance point $(0,M(s_{0}))$ exhibits a significant deviation from the regression line in both plots.
  • Intercept and slope are dependent if we fix the LLM and dataset. The general pattern is that when the intercept increases, the slope decreases. In fact, when the memorization score e= 50. Besides, $M(s_{50})\approx 11$. Although different regression lines have different slopes and intercepts, the y-value reaches about 11. This might indicate that intercept and slope may have a degree of freedom 1.
  • With a fixed temperature, enabling top-k or top-p sampling increases intercept and decreases slope.
  • The estimated normalized entropy decreases with the memorization score increasing.
The first two observations are consistent with observation points 2 and 3 in Section 5.
Table 4: Entropy-Memorization Law under different LLM sampling strategy
Strategy r Intercept Slope
Temp=0 0.933 5.490 0.106
Temp=0.5 0.936 5.474 0.106
Temp=0.8 0.926 5.011 0.113
Temp=0.8, top_p=0.5 0.935 5.599 0.103
Temp=1 0.944 4.646 0.118
temp=0.8, top_k=10 0.944 5.138 0.111
Figure 5: Temp=0
Figure 6: Temp=0.5
Figure 7: Temp=0.8
Figure 8: Temp=0.8, Top p=0.5
Figure 9: Temp=1
Figure 10: Temp=0.8, top-k=10
C.2 Entropy-Memorization Law Under Varying Generation Tokens Lengths
In the main body, we fixed the generated token length to 50. This subsection then explores different experimental setups to length $\{10,20,30,40,50\}$ generated tokens, respectively. We use OLMo-2-1124-7B and its training dataset OLMo-2-1124-Mix in this experiment. Note that we rescaled the memorization score to the range $[0,1]$ in the plot for better presentations.
Figure 11: Entropy-Memorization Law under varying generation tokens lengths.
Figures 11 present the experimental result when generation token length varies. Pearson correlation coefficients ($r$) range from 0.92 to 0.98 across all settings, indicating robust fitness under varying token lengths.
Another observation is that there is a monotonic increase in the intercept values of the fitted regression lines as the generation token length increases. This matches our expectations derived from information theory. Denote the vocabulary size as $|V|$, $n$-sequence has $|V|^{n}$ potential outcomes; hence when $n$ gets larger, the maximum entropy of $n$-sequence increases.
Figure 12: Estimated normalized entropy vs memorization score.
As a follow-up of section 6 in the main body, Fig. 12 presents the (estimated) normalized entropy v.s. memorization score. The findings are consistent with the findings that we observed in section 6. Moreover, $|r|$ increases from 0.82 to 0.97 as generation length decreases.
C.3 Entropy-Memorization Law Under Disparate Semantic Data
We employed a semantic-agonistic strategy to sample the dataset in the main body. This section then explores Entropy-Memorization Law under different semantic data. We chunk the sampled dataset into $k=16$ semantic clusters, develop a strategy to find the semantics of each cluster, and then examine EM Law under these disparate semantic data. This experiment was based on the OLMo-1B model using 240,000 training samples.
C.3.1 Semantic Clustering Pipeline
The specific steps are as follows:
  1. 1.
    Extracting semantic embedding. We first encode each answer sequence into embeddings employed using Sentence Transformers in Huggingfaces.
  2. 2.
    Clustering. With semantic embeddings, we apply K-Means [66] to partition the data into $k=16$ semantic clusters.
  3. 3.
    Identifying semantics of the cluster. We develop a pipeline that largely automates identifying the semantics of each cluster. The details will be shown in Appendix C.3.3.
  4. 4.
    Run Algorithm 2 on 16 partitions of the dataset. For each cluster, a linear regression is applied. We report the Pearson correlation coefficient, slope, and intercept and visualize the fitted lines.
For step 1, we select a popular pre-trained model all-mpnet-base-v2 [67] as the Sentence Transformer encoder to project a sentence to a high-dimension embedding. All-mpnet-base-v2 is built upon the MPNet architecture proposed by Microsoft and demonstrates superior performance across multiple semantic embedding benchmarks.
C.3.2 Clustering Result
Figure 13: Clustering Visualization (OLMo-1B Pretraining Datasets)
To visualize the clustering result, T-SNE is applied to reduce the dimensionality of the semantic embeddings to two dimensions. Figure 13 shows the result.
(a) Cluster 0
(b) Cluster 1
(c) Cluster 2
(d) Cluster 3
(e) Cluster 4
(f) Cluster 5
(g) Cluster 6
(h) Cluster 7
Figure 14: Clusters 0–7.
(a) Cluster 8
(b) Cluster 9
(c) Cluster 10
(d) Cluster 11
(e) Cluster 12
(f) Cluster 13
(g) Cluster 14
(h) Cluster 15
Figure 15: Clusters 8–15.
Figure 14 illustrates the Entropy-Memorization Law under disparate semantic data. Each cluster exhibits a clear linear pattern, with Pearson correlation coefficients above 0.75.
In general, different clusters exhibit distinctive entropy. For example, cluster 1 (Address and Organizational Identifiers) exhibits low intercept, while cluster 3 (Community Engagement and Activities) and 14 (Programming) exhibit high intercepts.
C.3.3 Interpreting Semantics of Each Cluster
To identify the semantics of each cluster, we build a pipeline that automates a large part of the manual efforts needed. The pipeline is as follows:
  1. 1.
    Detect distinctive samples within each cluster. Zhang et al. [68] formulates this task as a differential clustering problem and proposes a FINC method. To quantitatively measure semantic distinctions among the 16 clusters obtained via K-means, we conducted 16 FINC comparisons. For each cluster $C_{i}$, we set $C_{i}$ as the novel dataset and the union of the remaining 15 clusters as the reference set. The input to FINC is the sentence embeddings of all instances in the set, and FINC suggests the distinctive samples in the cluster.
  2. 2.
    Keywords summarization. In this stage, we use tri-grams as effective descriptors for naming and interpreting cluster identities. Specifically, we use i) spaCy [69] to perform named entity recognition and dependency parsing to ensure that extracted units are linguistically complete phrases (e.g., “protective spell harry”, “lend broom fly”), and ii) YAKE [70] to ranks terms using heuristics such as frequency, context, and positional distribution.
  3. 3.
    Human annotation. Based on summarized keywords, human annotators further summarize the semantics of the cluster.
Semantics of each cluster
Table 5 presents top-5 keywords and human-annotated semantic labels for each cluster.
Table 5: Semantics of each cluster (cluster 0-7)
Cluster 0: Scientific Research and Technology Score Keyword Phrase 1.28e-04 translate depth direction 1.28e-04 design fabrication characterization 1.28e-04 pct design fabrication 1.28e-04 manipulation pct design 1.23e-04 sensor control manipulation
Cluster 1: Address and Organizational Identifiers Score Keyword Phrase 1.74e-05 street number city 1.54e-05 party committee number 1.54e-05 city number 1.52e-05 type number form 1.46e-05 conduit state number
Cluster 2: Religious and Biblical Texts Score Keyword Phrase 7.94e-05 people sword thy 7.94e-05 thy people sword 7.91e-05 jacob son reuban 7.84e-05 son brother house 7.80e-05 son lord hath
Cluster 3: Community Engagement and Activities Score Keyword Phrase 1.05e-04 irrigation evaporate leave 1.05e-04 bullet time jump 1.04e-04 community meetup world 1.03e-04 community kid spout 1.02e-04 year electronic music
Cluster 4: Social Development and Policy Score Keyword Phrase 1.34e-04 dramatically drive price 1.33e-04 develop skill team 1.33e-04 prefer policy influence 1.33e-04 outwith prefer policy 1.33e-04 sector real passion
Cluster 5: Software Development Configuration Score Keyword Phrase 2.39e-05 true plugin proposal 2.29e-05 node optional true 2.27e-05 header content type 2.26e-05 true ellipsis true 2.26e-05 dev true child
Cluster 6: Spiritual and Religious Beliefs Score Keyword Phrase 9.67e-05 thy mind thy 9.61e-05 death eternal life 9.54e-05 create sustain universe 9.52e-05 world drive ulterior 9.52e-05 behavior world drive
Cluster 7: Health Risks and Medical Studies Score Keyword Phrase 1.06e-04 health increase risk 1.06e-04 evaluate risk factor 1.05e-04 pregnancy increase risk 1.05e-04 disease high risk 1.05e-04 high disease risk
Table 6: Semantics of each cluster (cluster 8-15)
Cluster 8: Economic Indicators and Growth Score Keyword Phrase 2.54e-05 billion country oda 2.51e-05 permanent surface runway 2.51e-05 kwh capita growth 2.50e-05 imf intelsat interpol 2.48e-05 price growth rate
Cluster 9: Evolutionary Biology and History Score Keyword Phrase 4.49e-04 primatology million year 4.48e-04 mutate gene ancient 4.48e-04 start million million 4.45e-04 paleolithic million year 4.44e-04 account million year
Cluster 10: Fantasy Narrative (Harry Potter) Score Keyword Phrase 4.98e-05 dumbledore harry meet 4.95e-05 forward harry ron 4.94e-05 muggle bystander incredulously 4.72e-05 sirius black peril 4.71e-05 harbor end time
Cluster 11: Personal Experiences and Emotions Score Keyword Phrase 1.07e-04 hunt season roll 1.06e-04 love good love 1.06e-04 thing work happen 1.05e-04 decision normal result 1.04e-04 understand thing happen
Cluster 12: Linguistic and Semantic Analysis Score Keyword Phrase 1.22e-04 interpret apply male 1.20e-04 ingen det finne 1.20e-04 finne ingen det 1.19e-04 title aktivt medlem 1.17e-04 confusion attain agenda
Cluster 13: Genetic and Biomedical Research Score Keyword Phrase 8.51e-05 mesenchymal stem cell 8.40e-05 screen gene foster 8.27e-05 interaction show high 8.26e-05 expression level gene 8.25e-05 search perform blast
Cluster 14: Programming Command Scripts Score Keyword Phrase 4.64e-06 text text ohm 4.57e-06 delaytimer command 4.03e-06 dark text text 3.94e-06 command runcmd 3.66e-06 cost command type
Cluster 15: Political Events and Commentary Score Keyword Phrase 1.41e-04 white house official 1.41e-04 idea african americans 1.38e-04 african americans woman 1.38e-04 end war year 1.37e-04 snc lavalin scandal
Appendix D A Detailed Example of Gibberish Memorization
Prompt Text:
1==} engines: {node: ^10.0.0 $|$ ^12.0.0 $|$ >= 14.0.0} peerDependencies: eslint: ’>=5’ dependencies: eslint: 8.5.0 eslint-visitor-keys: 2.1.0 dev: true /eslint-visitor-keys/1.3.0: resolution: {integrity: sha512-6J
Generated Text (exactly match with answer text):
172N8UNa462wa/KFODt/PJ3IU60SDpC3QXC1Hjc1BXXpfL2C9R5+AU7jhe0F6GREqVMh
This prompt is part of a configuration file. Following the prefix of a SHA-512 value, OLMo-1B perfectly memorizes the value after the prefix.
Appendix E Extended Results on Dataset Inference
E.1 Entropy Memorization Law on Test Data
This subsection presents the plot of applying the entropy estimator and normalized entropy estimator to test data. To maintain consistency throughout the paper, we use the terms used for memorization with slight abuse. For example, the memorization score still measures the distance between the ground truth and the model’s response, but the model is not actually "memorizing" training data. Following the same setup described in the main body, we select data from LiveBench [39] from 2024-06-25 to 2024-11-25. The time property guarantees that LiveBench is non-member data for OLMo-2-1124-7B. LiveBench is licensed under CC BY-SA 4.0 International License.
Figure 16: Entropy Memorization Law on OLMo-2 with LiveBench dataset.
With around 3k samples from LiveBench, we observe that:
  • Level-set-based entropy is still a good indicator of LLM performance, as demonstrated by the high Person’s $r$. The sample space also approximately grows exponentially as the memorization score increases.
  • However, compared with training data, the intercept is lower, and the slope is higher. There is less coverage of low memorization scores, especially for memorization scores within 0-10.
  • at low-distance set, it is observed that OLMo-2 is producing low-entropy text. The entropy is much lower than what we observed on train data.
Here is a case study on a 2-distance memorization:
Prompt Text:
[…]
Generated Text (which exactly match with answer text):
$\n1.0,0.0,0.0,0.0,0.0$$0.0\n$2.0
Given the prompt given in this case study, OLMo-2 generates repeated numbers, exhibiting low entropy. This observation further inspires EMBEDI to perform a dataset inference task, which we will cover later.
E.2 EMBEDI: Entropy-Memorization law-BasEd Dataset Inference
EMBEDI is inspired by several empirical observations. First, over-parametrization of LLMs may lead to overfitting on training data, resulting in a generalization gap between training data and testing data; then, LLM may perform well on low-entropy testing data, resulting in a low intercept in EM Law. Besides, given a fixed dataset and LLM, empirical evidence suggests that the intercept and slope generated by Algorithm 2 are dependent. This inspires us to develop the following strategy for dataset inference:
Given an LLM $\theta$ and dataset $D$, run Algorithm 2 and get intercept $k$. Compare $k$ with a pre-defined threshold $\tau_{k}$. Assign label 1 (i.e., member) if $k>\tau_{k}$, assign label 0 otherwise.
Note that in the main body, we present a variant of the strategy above; however, based on our observation that "intercept and slope are dependent," two variants should be equivalent.
Amount of data required for EMBEDI
Effective dataset inference requires reliable entropy estimation and diverse memorization score distributions. In section 6, we have revealed that the frequency of a low memorization score is exponentially smaller than that of a high memorization one. We, therefore, set the minimum sample size to n=1,500n=1{,}500, with each sample is a 150-token sequence.
E.3 Experimental Results on Dataset Inference
Table 7: Extended result on dataset inference
LLM Dataset Intercept Slope Prediction Label
OLMo-2 LiveBench 2.202 0.155 0 0
Pythia MIMIR_cc -2.048 0.251 0 0
Pythia MIMIR_cc 3.992 0.091 1 1
OLMo-2 OLMo-2-1124-Mix 3.724 0.142 1 1
Pythia MIMIR_full 6.297 0.092 1 0
Pythia MIMIR_full 6.166 0.095 1 1
Pythia MIMIR_tarxiv 1.156 0.174 1 1
Pythia MIMIR_tarxiv -0.910 0.0227 0 0
Pythia MIMIR_wiki 3.006 0.131 1 0
Pythia MIMIR_wiki 2.894 0.133 1 1
Selected LLMs. We select OLMo-2-1124-7B (“OLMo-2”), Pythia-6.9b-deduped [16] (“Pythia”).
Selected Datasets. We select LiveBench [39] and MIMIR [40]. For LiveBench, we use data from 2024-06-25 to 2024-11-25. MIMIR is a public dataset originally for evaluating MIAs on the Pythia suite by re-compiling the Pile [41] train/test splits. For MIMIR, we use “Pile CC” “temporal arXiv”, “wiki” subset, and full dataset (“full”) for evaluation. Note that LiveBench and Temporal arxiv are temporal-cutoff-based, while the remaining dataset is i.i.d.-based.
Threshold for each LLM. In EMBEDI, we assign a threshold for each LLM. We empirically set $\tau_{k}$ to 0 and 3 for Pythia and OLMo-2, respectively. Other variants of thresholding strategy include domain-specific thresholding, and we leave it to future work.
Table 7 presents the extended results of EMBEDI on the dataset inference task.
Discussions
EMBEDI enjoys several advantages: it is compute-efficient – it only requires LLM inference on $n$ samples. No training is required. It does not require any additional reference models.
However, we regard that the methodology of EMBEDI is fully based on empirical observations, the theoretical grounds behind remain unclear.